∫(x^-11)/Sqrt(1+x^4) dx Is equal to how much_------------@------------
                           

Grade:12th pass

1 Answers

Subramanyam
5 years ago
∫1/(x^13*(sqrt(1+1/(x^4))))dx---(4)
1+1/(x^4)=t^2----(1)
1/(x^8)=(t^2-1)^2---(2)
-4/(x^5)dx=2tdt----(3)substitute these 3 in (4)